Niceshotbroski:

A popular new pair of shoes are sold at the shoe store for a regular price of $85. The shoe store currently has a sale of 20% off the original price of all new shoes. The tax rate at the shoe store is 7% and is charged based on the discounted price of the shoes.

Niceshotbroski:

anyone?

GeekGold:

So the answer should be the original price times .20 (for the discount) and than add the tax by multiplying the discounted price by .07 and adding that to the discounted price

Niceshotbroski:

Oh ok

GeekGold:

Here's the equation that should be used. f(x)=(x-(x*.20))+(x*.07)

GeekGold:

I believe

Niceshotbroski:

Ok

Extrinix:

So, to find the amount off of the original price, multiply 85 by 0.20 \(85\times 0.20= 17\) Then, reduce the price of 85 by 17 \(85-17=68\) After that you want to find the tax rate of 68 \(68 \times 0.07=4.76\) Now you want to add the tax amount to 68 \(68+4.76=72.76\) So the price with discount and tax would be `72.76`

Extrinix:

@geekgold wrote:
I believe
Oops, didnt see that x.x

Niceshotbroski:

Thanks :)

GeekGold:

It's fine, you explained it a lot better

GeekGold:

XD

Extrinix:

No problem ^^ (from both of us, not just me x.x)

Niceshotbroski:

thanks im doing a mid unit rn